OpenStudy (shubhamsrg):

finding 100th digit after the decimal place of ( 2 + sqrt5)^500

OpenStudy (shubhamsrg):

I do have the solution but I do not understand it. In case anyone would like to see the solution ?

OpenStudy (agent0smith):

Hmm, I'm guessing maybe we need to use the binomial theorem... all the even powers on the sqrt 5 would give us whole numbers, all the odd powers would give multiples of sqrt 5.

OpenStudy (anonymous):

I would like to see the solution, when you are online, post it, please.

OpenStudy (shubhamsrg):

http://yourmathguru.com/forums/topic/the-100th-digit

OpenStudy (anonymous):

@mukushla look here if you can help then..

OpenStudy (shubhamsrg):

are you sure wolfram is telling 100th digit after the decimal ?

OpenStudy (shubhamsrg):

well yes but what if you want to know 3rd digit after decimal of 186.3349 ?? input will vary

OpenStudy (phi):

yes, wolfram is a tool. Maybe in a few more years it will be a mathematician ...

OpenStudy (shubhamsrg):

well what I wanted to say was your input in wolfram was wrong

OpenStudy (phi):

oh, wait. It is the 10^313 power, so that is not what you want.

OpenStudy (phi):

so here is what you want sage: N(( 2 + sqrt(5))^500,digits=420) 3.03012381665493808617352705454714235391879140029371493970855922136019868971187806629576822563956581117434290608358832071471866334615305238079729662622640319590069101303430750653363699328946679390632157734812572428455630808247147382467591044592551304368531540021043660638089720375493820280432347278412694880050000019999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999998e313 (the last 8 may not be accurate) so it is 9.

OpenStudy (phi):

I was just curious if we could get the answer using computation rather than analysis.

OpenStudy (shubhamsrg):

hmm.. well since we are confirmed answer is 9 , did you understand their method ?

OpenStudy (phi):

Here is what they did. (2+ sqrt5)^500 + (2 - sqrt5)^500 is an integer use the binomial theorem to expand and note that you get terms like nCm 2^m 5^{(n-m)/2} + nCm (-1)^(n-m) 2^m 5^{(n-m)/2} when n-m is odd, these cancel. when n-m is even you get terms nCm 2^(m+1) 5^{(n-m)/2} which are integers. next, we have 2-sqrt5 is about -0.236..... raised to the 500th power (2-sqrt5)^500 will be positive and much less than 1 0 < (2-sqrt5)^500 << 1 we can take the log base 10 of (2-sqrt5)^500 to get an estimate of big (small ) this number is. 500*log(0.236...) (we use a positive base to avoid imaginary numbers) we get about -313.5 in other words, (2-sqrt5)^500 about = 10^-313.5 it is 0.0000... lots of zeros ...00000...digits on the other hand, (2+sqrt5)^500 is some big number nnnn.dddddd with an integer part and a fractional part when we add (2-sqrt5)^500 to it, we get an integer (no decimals) nnnnn...nnn.ddddddddddddddddddddddd +000000000.000000000000000000eeeeee ___________________________________________ mmmmmm.000000000000000000000000 (btw, d is not necessarily the same digit here) we know the eee part must add with the ddd part to give zeros, which means we are generating a carry of 1 at the 313 digit in we have d + 0 + carry= 0 or because we know we have a carry of 1 d+1=0 plus carry d must be 9, which gives 0 plus a carry. Following that logic, all the d's in the 312 position or less must be 9, to generate 0's therefore the 100th digit must be a 9

OpenStudy (phi):

still not clear ?

OpenStudy (shubhamsrg):

I understood the log thing . but it bothers me that in this question you'll have to make use of calculator